OpenStudy (anonymous):

s=vt−1/2at2 ; for t.

OpenStudy (anonymous):

Isolate t?

OpenStudy (anonymous):

\[0 = -\dfrac{1}{2}at^2 + vt - s\] Use quadratic formula: A = \(-\dfrac{1}{2}a\); B = v; C = -s \[t = \dfrac{-B \pm \sqrt{B^2 - 4AC}}{2A}\]

OpenStudy (anonymous):

i tried but i didnt get the right answer, please help me solve ill give u a medal n become ur fan!!